Tutors please help me: THANK YOU
solve the exponential equation 3^x2+^4x=1/27
and

28^x=10^-3x

Answer by stanbon(48569)

solve the exponential equation 3^(x^2+4x) = 1/27
3^(x^2+4x) = 3^-3
x^2+4x=-3
x^2+4x+3=0
(x+3)(x+1) = 0
x = -3 or x=-1
Checking those possible answers:
x= -3 check out
x= -1 does not
------------------------
28^x=10^-3x
xlog28 = -3x
3x+xlog28 = 0
x(3+log28) = 0
x = 0
====================
Cheers,
Stan H.